OpenStudy (cherie_magee):

which of the following ordered pairs is solution of x+1/2y=1?

OpenStudy (anonymous):

which ordered pair?

OpenStudy (cherie_magee):

(-2,6) (2,-6) (-2,_6) @FriedRice

OpenStudy (anonymous):

just plug in the points into the equation

OpenStudy (anonymous):

so first answer -2+1/2(6)=1 -1/12 = 1 so no

OpenStudy (anonymous):

second question2+ 1/2(-6) = 1 3/-12 =1 so no

OpenStudy (anonymous):

the third -2 + 1 / 2( -6) = 1 1/ -12 = 1 so no

OpenStudy (nurali):

x+1/2y=1 the ordered pair i will use is the first one. (-2, 6) so (-2) + 1/2(6)=1 1/2(6)=3 so (-2) + 3 = 1 yes this ordered pair works for the solution meaning it is the answer.

OpenStudy (anonymous):

ohhh I didn't get the question stupid me

OpenStudy (cherie_magee):

-_- @FriedRice

OpenStudy (anonymous):

lol